How many different groups can be selected for playing tennis out of 4 ladies and 3 gentlemen,there being one lady and one gentlemen on each side

Answer by Edwin McCravy(20055)

Each foursome consists of two ladies and 2 gentlemen.

We select the ladies 

4 ladies choose 2 = 4C2

and

3 gentlemen choose 2 = 3C2

Answer: 4C2×3C2 = 6×3 = 18

Edwin